Find the equation of the curve at every point that passes through the point (0, −1) and which the normal line at any point (x, y) has a slope of 1/(3x + 2).

1
Expert's answer
2021-12-21T17:01:18-0500

Let the equation of the curve be "y=f(x)." Then slope of the nornal to the curve at any point "(x, y)" will be

"slope=m=-\dfrac{1}{f'(x)}"

Given "slope=m=\dfrac{1}{3x+2}."

Then


"f'(x)=-\dfrac{1}{m}=-(3x+2)"

Integrate


"f(x)=-\int(3x+2)dx=-\dfrac{3}{2}x^2-2x+C"

The curve passes through the point "(0, \u22121)"


"-1=-\dfrac{3}{2}(0)^2-2(0)+C=>C=-1"

The equation of the curve at every point is


"y=-\dfrac{3}{2}x^2-2x-1"